# MTVQA: Benchmarking Multilingual Text-Centric Visual Question Answering

### v1

@@ Abstract
Text-Centric Visual Question Answering (TEC-VQA) in its proper format not only facilitates human-machine interaction in text-centric visual environments but also serves as a de facto gold proxy to evaluate AI models in the domain of text-centric scene understanding. However, most TEC-VQA benchmarks have focused on high-resource languages like English and Chinese. Despite pioneering works to expand multilingual QA pairs in non-text-centric VQA datasets using translation engines, the translation-based protocol encounters a substantial “Visual-textual misalignment” problem when applied to TEC-VQA. Specifically, it prioritizes the text in question-answer pairs while disregarding the visual text present in images. Furthermore, it does not adequately tackle challenges related to nuanced meaning, contextual distortion, language bias, and question-type diversity. In this work, we address the task of multilingual TEC-VQA and provide a benchmark with high-quality human expert annotations in 9 diverse languages, called MTVQA. To our knowledge, MTVQA is the first multilingual TEC-VQA benchmark to provide human expert annotations for text-centric scenarios. Further, by evaluating several state-of-the-art Multimodal Large Language Models (MLLMs), including GPT-4V, on our MTVQA dataset, it is evident that there is still room for performance improvement, underscoring the value of our dataset. We hope this dataset will provide researchers with fresh perspectives and inspiration within the community.
footnotetext: Equal contribution. 🖂Corresponding author.
@@ 1 Introduction
In the era of burgeoning AI, especially in LLMs/MLLMs [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17], Text-Centric Visual Question Answering (TEC-VQA[18, 19, 10, 9, 13, 12, 11] has served as a de facto gold proxy to evaluate AI models in the domain of text-centric scene understanding. Compared with general VQA [18, 20, 21, 19, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33], TEC-VQA places greater emphasis on answering questions that require understanding textual information within images. It provides a streamlined avenue for individuals without specialized expertise to articulate their requirements and access applications in text-centric visual environments. However, the majority of advancements in TEC-VQA have predominantly concentrated on high-resource languages, e.g., English [18, 19, 20, 23], Chinese [34, 31], Japanese [35, 36] and etc., thus restricting the applicability of AI models to the global community, particularly populations speaking low-resource languages.
To tackle the problem of language diversity, several seminal studies [37, 38, 39] in the general VQA field, leverage off-the-shelf translation engines to expand existing question-answer pairs from high-resource languages to their multilingual counterparts including low-resource ones. However, when applied to TEC-VQA, this translation-based approach may fall prey to the “Visual-textual misalignment” problem as only the text in question-answer pairs can be processed, while the visual text present in the images is overlooked. Not to mention issues such as nuanced meaning, contextual distortion, language bias, and question type diversity further render the transferability of the translation protocol infeasible for TEC-VQA. The status quo begs for a question: “How can we address the visual-textual misalignment problem for multilingual TEC-VQA and what we stand in the MLLM era?
@@
Figure 1: Left: overview of various categories of text-centric images. Right: distribution of image and QA pairs over the 9 languages in MTVQA benchmark.
Figure 2: Multilingual VQA examples selected from four languages. From left to right: Arabic (AR), Russian (RU), Thai (TH), Vietnamese (VI). The corresponding translations in English are in brackets.
In this work, to answer the question above, we establish MTVQA, a novel and high-quality multilingual TEC-VQA benchmark, where all images are collected from real-world and meticulously annotated by human experts in nine languages: Arabic (AR), Korean (KO), Japanese (JA), Thai (TH), Vietnamese (VI), Russian (RU), French (FR), German (DE), and Italian (IT). More concretely, to ensure the visual-textual alignment at most, the annotation process follows the raise-then-correct paradigm, where a group of human annotators raises several distinct questions, ranging from simple content extraction to text-related reasoning, and subsequently provides answers. These QA pairs are then double-checked by another group to ensure accuracy and consistency. Consequently, as illustrated in Fig. 5, 6,678 training images and 21,829 question-answer pairs, as well as 2,116 test images and 6,778 question-answer pairs are obtained, covering several fine-grained scenarios, such as menus, logos, maps, bills, PPTs, research papers, and etc. To our best knowledge, MTVQA is the first TEC-VQA dataset to provide native human annotations for multilingual text-rich scenarios, especially for low-source languages. Furthermore, we investigate recent representative MLLMs, including GPT-4V, Gemini, QwenVL etc., by juxtaposing experimental results regarding their performance on our newly proposed MTVQA. Both for general MLLMs and document-focused ones, the results unequivocally demonstrate that opportunities for improvement persist within these MLLMs when applied in multilingual text-rich scenarios.
In summary, the main contributions of this paper can be categorized into three points:
  • We introduce the MTVQA dataset, to the best of our knowledge, which is the first multilingual TEC-VQA benchmark to provide human expert annotations for text-centric scenarios.
  • We benchmark the state-of-the-art MLLMs on our new dataset and show there is still room for performance improvement for these models under multilingual text-rich scenarios.
  • We propose a set of baselines for multilingual TEC-VQA tasks.
@@ 2 Related Work
2.1 LLMs/MLLMs for text-centric VQA
Recent advancements in LLMs/MLLMs [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17] have revolutionized VQA tasks, as demonstrated by the remarkable zero-shot performance of these models. Notably, the high generalizability of LLMs/MLLMs, when explicitly trained on visual text understanding datasets and fine-tuned with instructions, has significantly enhanced their application in text-centric VQA scenarios [10, 9, 13, 12, 11]. For example, LLaVAR [40], UniDoc [10], which extend LLaVA [30] into the realm of document understanding, pioneering the text-centric VQA of MLLMs by training them to predict texts and coordinates from document images. Furthermore, DocPedia [9] operates visual input in the frequency domain rather than in space, which enables higher input resolution without increasing the input sequence. Lately, mPLUG-DocOwl [41], Qwen-VL [6], and TextMonkey [12] leverage publicly available document-related VQA datasets to further enhance the text-centric VQA capabilities. Despite the promising results achieved by existing LLMs/MLLMs in text-centric VQA tasks, their focus on high-resource languages such as English or Chinese has posed challenges in achieving reasonable performance for low-resource languages. This is primarily due to the lack of data or benchmarks for these low-resource languages.
@@ 2.2 Multilingual text-centric VQA Benchmarks
There is a recent effort toward extending VQA tasks to a wider range of languages [42, 38, 43, 39, 44, 37] by providing a multilingual VQA datasets. For example,  Gao et al. [31] created a free-form bilingual VQA dataset (FM-IQA) contains over 150,000 images and 310,000 freestyle Chinese question-answer pairs and their English translations. Raj Khan et al. [37] developed a large-scale multilingual and code-mixed VQA dataset (MuCo-VQA) supporting five languages. Of more relevance are the works xGQA (8 languages) [38] and MaXM (7 languages) [39], which apply translation-based protocols to expand VQA data beyond English. However, the translation-based multilingual VQA datasets inherently face issues, such as the “Visual-textual misalignment” problem, where only the text in question-answer pairs is processed, while the visual text in images is overlooked. Additionally, the nuanced meaning and context are often distorted; language bias introduced by machine translation models, and the coverage of certain question types is limited, as highlighted by Changpinyo et al. [39]. Moreover, none of the previous multilingual datasets focus on text-centric scenarios where multilingual text frequently occurs.
Our benchmark distinguishes itself by focusing on multilingual text-centric VQA scenarios using human expert annotations. To the best of our knowledge, the MTVQA benchmark is the first dataset to provide native human annotations for such scenarios. It covers 9 languages, thereby facilitating the training and evaluation of multilingual models in diverse linguistic contexts. Additionally, our dataset can gauge the VQA system’s ability for not only high-resource languages but also those that are typically underrepresented in current datasets [18, 19, 20, 23, 31].
The MTVQA benchmark addresses a significant gap in existing datasets by catering to the crucial needs of low-resource languages through annotations from native speakers across multiple languages. Our pioneering efforts distinctly position the MTVQA benchmark as a unique multilingual VQA resource, advancing the frontier of machine learning research.
@@ 3 MTVQA Benchmark
The MTVQA Benchmark covers 9 languages: Arabic (AR), Korean (KO), Japanese (JA), Thai (TH), Vietnamese (VI), Russian (RU), French (FR), German (DE), and Italian (IT). In this section, we describe in detail how we establish the MTVQA benchmark, including the collection of raw image data and two-round human expert annotations, which are independent of each other.
@@ 3.1 Data Collection
Our purpose is to develop a multilingual VQA benchmark capable of evaluating the QA performance of MLLMs in multilingual text-centric scenarios, thus the raw data collection process is mainly oriented towards text-centric images from natural scenarios and document scenarios. To ensure the diversity and quality of data, we collect not only the raw image data from publicly available datasets, including the multilingual scene text recognition images from MLT2019 [45] and PowerPoint slides (PPTs) sourced from the internet, but also the data from countries of each language. Furthermore, the collected data includes multiple fine-grained scenarios (Fig. 1), such as menus, logos, maps, bills, PPTs, research papers, and etc. As a result, we gather a total of 1,220 images from document scenarios and 876 images from natural scenarios in the test set of the MTVQA benchmark. To ensure the visual-textual alignment, for text-rich images lacking text and language annotations, we subject them to a standardized data cleaning process, which includes text recognition and language classification. Afterward, we organize all the text-rich images we have obtained into language-specific groups, preparing them for the subsequent stage of data annotation.
@@ 3.2 Human Expert Annotation
In order to obtain informative and accurate text-related QA pairs on the language-specific grouped images, we recruit a group of annotators with expertise from local regions of each language. It is worth noting that all these annotators are native speakers of their respective languages, ensuring their deep understanding and proficiency in the linguistic nuances and cultural context necessary for precise annotations. Considering the subjective nature of the text-image understanding task, we have implemented a further division within the annotation team. This division involves separating the team into two independent groups, with one group dedicated to generating and responding to questions based on the provided images, while the other group focuses on evaluating and correcting the QA pair results. This raise-then-correct paradigm ensures a comprehensive and reliable assessment of the text-image understanding process. Additionally, each language’s annotation results undergo a 10% sampling inspection by a quality inspector. If the QA pairs fail to meet the criteria, they are sent back for re-annotation. Prior to commencing the formal human expert annotation task, all annotators undergo unified training and receive annotation examples. The brief diagram of the two-round annotation process is shown in Figure 3 and we elaborate on it in the following subsections.

### v2

@@ Abstract
Text-Centric Visual Question Answering (TEC-VQA) in its proper format not only facilitates human-machine interaction in text-centric visual environments but also serves as a de facto gold proxy to evaluate AI models in the domain of text-centric scene understanding. Nonetheless, most existing TEC-VQA benchmarks have focused on high-resource languages like English and Chinese. Despite pioneering works to expand multilingual QA pairs in non-text-centric VQA datasets through translation engines, the translation-based protocol encounters a substantial “visual-textual misalignment” problem when applied to TEC-VQA. Specifically, it prioritizes the text in question-answer pairs while disregarding the visual text present in images. Moreover, it fails to address complexities related to nuanced meaning, contextual distortion, language bias, and question-type diversity. In this work, we tackle multilingual TEC-VQA by introducing MTVQA, the first benchmark featuring high-quality human expert annotations across 9 diverse languages, consisting of 6,778 question-answer pairs across 2,116 images. Further, by comprehensively evaluating numerous state-of-the-art Multimodal Large Language Models (MLLMs), including GPT-4o, GPT-4V, Claude3, and Gemini, on the MTVQA dataset, it is evident that there is still a large room for performance improvement, underscoring the value of MTVQA. Additionally, we supply multilingual training data within the MTVQA dataset, demonstrating that straightforward fine-tuning with this data can substantially enhance multilingual TEC-VQA performance. We aspire that MTVQA will offer the research community fresh insights and stimulate further exploration in multilingual visual text comprehension. The project homepage is at https://bytedance.github.io/MTVQA/.
footnotetext: Equal contribution. 🖂 Corresponding authors.
@@ 1 Introduction
In the era of burgeoning AI, especially in MLLMs [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18], Text-Centric Visual Question Answering (TEC-VQA[19, 20, 11, 10, 14, 13, 12] has served as a de facto gold proxy to evaluate AI models in the domain of text-centric scene understanding. Compared with general VQA [19, 21, 22, 20, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34], TEC-VQA places greater emphasis on answering questions that require understanding textual information within images. It provides a streamlined avenue for individuals without specialized expertise to articulate their requirements and access applications in text-centric visual environments. However, the majority of advancements in TEC-VQA have predominantly concentrated on high-resource languages, e.g., English [19, 20, 21, 24], Chinese [35, 32], Japanese [36, 37] and etc., thus restricting the applicability of AI models to the global community, particularly populations speaking low-resource languages.
To tackle the problem of language diversity, several seminal studies [38, 39, 40] in the general VQA field, leverage off-the-shelf translation engines to expand existing question-answer pairs from high-resource languages to their multilingual counterparts including low-resource ones. However, when applied to TEC-VQA, this translation-based approach may fall prey to the “visual-textual misalignment” problem as only the text in question-answer pairs can be processed, while the visual text present in the images is overlooked. Although the visual text can be obtained by an OCR (Optical Character Recognition) engine and then translated into English, this indirect process could introduce issues such as nuanced meaning, contextual distortion, language bias, and question type diversity. These factors further render the transferability of the translation protocol infeasible for TEC-VQA. Taking the second case in Fig. 1 as an example, if either the visual text or the question in Russian is recognized or translated problematically, then the question would never be answered. The status quo begs for a question: “Can we address the visual-textual misalignment problem directly from the source language per se for multilingual TEC-VQA and what we stand in the MLLM era?
Figure 1: Multilingual text-centric VQA visualization selected from four languages. From left to right: Arabic (AR), Russian (RU), Thai (TH), Vietnamese (VI). The corresponding translations in English are in brackets. More examples see Figure 7.
@@
Figure 2: Left: overview of various categories of text-rich images. Right: image and QA pairs distribution over the 9 languages in MTVQA benchmark.
In this work, to answer the question above, we establish MTVQA, a novel and high-quality multilingual TEC-VQA benchmark, where all images are collected from real-world and meticulously annotated by human experts in nine languages: Arabic (AR), Korean (KO), Japanese (JA), Thai (TH), Vietnamese (VI), Russian (RU), French (FR), German (DE), and Italian (IT). More concretely, to ensure the visual-textual alignment at most, the annotation process follows the raise-then-correct paradigm, where a group of human annotators raises several distinct questions, ranging from simple content extraction to text-related reasoning, and subsequently provides answers. Another group then double-checks these QA pairs to ensure accuracy and consistency. Consequently, as illustrated in Fig. 2, 6,678 training images and 21,829 question-answer pairs, as well as 2,116 test images and 6,778 question-answer pairs are obtained, covering more than 20 fine-grained scenarios from both documents and natural scenes, such as menus, logos, maps, bills, PPTs, research papers, and etc. To our knowledge, MTVQA is the first TEC-VQA dataset to provide native human annotations for multilingual text-rich scenarios, especially for low-source languages.
Figure 3: Left: comparison of the overall performance of various MLLMs in the MTVQA benchmark. Right: comparison of the performance exhibited by MLLMs in the 9 languages of the MTVQA.
We further investigate recent representative MLLMs in Fig. 3, including GPT-4o, GPT-4V, Gemini, QwenVL, etc., and their performance on our proposed MTVQA. As vividly revealed in Fig. 3 (a), GPT-4o is the top performer, followed by Claude3 Opus and Gemini Ultra, with closed-source MLLMs generally leading. General MLLMs like QwenVL Max and QwenVL Plus show mid-tier performance, while text-centric MLLMs lag behind. Fig. 3(b) shows closed-source MLLMs consistently outperforming others, especially in French (FR) and German (DE), while languages like Arabic (AR) and Thai (TH) witness lower scores across all categories. The results unequivocally demonstrate that opportunities for improvement persist within existing MLLMs when applied in multilingual text-rich scenarios. In summary, the main contributions of this paper can be categorized into three points:
  • We coin the MTVQA dataset, which to the best of our knowledge is the first multilingual TEC-VQA benchmark providing human expert annotations to solve “visual-textual misalignment” problem in multilingual text-centric scenarios.
  • We benchmark the state-of-the-art MLLMs on our new dataset and show that there are still opportunities for improvement on even the most advanced MLLMs when it comes to multilingual text-rich scenarios.
  • We establish a set of new multilingual TEC-VQA baselines for both closed-source and general-purpose, text-centric MLLMs.
@@ 2 Related Work
2.1 MLLMs for Text-centric VQA
Recent advancements in MLLMs [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18] have revolutionized VQA tasks, as demonstrated by the remarkable zero-shot performance of these models. Notably, the high generalizability of MLLMs, when explicitly trained on visual text understanding datasets and fine-tuned with instructions, has significantly enhanced their application in text-centric VQA scenarios [11, 10, 14, 13, 12]. For example, LLaVAR [41], UniDoc [11], which extend LLaVA [31] into the realm of document understanding, pioneering the text-centric VQA of MLLMs by training them to predict texts and coordinates from document images. Furthermore, DocPedia [10] operates visual input in the frequency domain rather than in space, which enables higher input resolution without increasing the input sequence. Lately, mPLUG-DocOwl [42], Qwen-VL [7], and TextMonkey [13] leverage publicly available document-related VQA datasets to further enhance the text-centric VQA capabilities. Despite the promising results achieved by existing MLLMs in text-centric VQA tasks, their focus on high-resource languages such as English and Chinese has posed challenges in achieving reasonable performance for low-resource languages. This is primarily due to the lack of data or benchmarks for these low-resource languages.
@@ 2.2 Multilingual Text-centric VQA Benchmarks
There are some recent efforts toward extending VQA tasks to a wider range of languages [43, 39, 44, 40, 45, 38] by providing a multilingual VQA datasets. For example,  Gao et al. [32] created a free-form bilingual VQA dataset (FM-IQA) containing over 150,000 images and 310,000 freestyle Chinese question-answer pairs and English translations. Raj Khan et al. [38] developed a large-scale multilingual and code-mixed VQA dataset (MuCo-VQA) supporting five languages. Of more relevance are the works xGQA (8 languages) [39] and MaXM (7 languages) [40], which apply translation-based protocols to expand VQA data beyond English. However, the translation-based multilingual VQA datasets inherently face issues, such as the “visual-textual misalignment” problem, where only the textual information in question-answer pairs is considered, while the visual text in images is overlooked. Additionally, the nuanced meaning and context are often distorted; language bias is introduced by machine translation models, and the coverage of certain question types is limited, as highlighted by [40]. Moreover, none of the previous multilingual datasets focus on text-centric scenarios where multilingual text frequently occurs.
Our benchmark, MTVQA, distinguishes itself by focusing on multilingual text-centric VQA scenarios using human expert annotations. It covers 9 languages, thereby facilitating the training and evaluation of multilingual models in diverse linguistic contexts. Additionally, our dataset can gauge the VQA system’s ability for not only high-resource languages but also those that are typically underrepresented in current datasets [19, 20, 21, 24, 32].
@@ 3 MTVQA Benchmark
The MTVQA benchmark is meticulously established to evaluate the multilingual text comprehension performance of Multimodal Large Language Models (MLLMs). MTVQA covers nine languages: Arabic (AR), Korean (KO), Japanese (JA), Thai (TH), Vietnamese (VI), Russian (RU), French (FR), German (DE), and Italian (IT). The data construction process involves a comprehensive collection of text-rich images and a two-round human expert annotation process.
Figure 4: A brief diagram of the annotation process.
@@ 3.1 Data Collection
The raw data collection is oriented towards gathering text-rich images from various natural and document scenarios, ensuring diversity and quality. This includes images from publicly available datasets(e.g., ICDAR MLT19 [46]) and those sourced from the internet, such as menus, logos, maps, bills, PowerPoint slides (PPTs), research papers, etc, as shown in Fig. 2. A total of 1,220 images from document scenarios and 876 images from natural scenarios are collected for the test set of the MTVQA benchmark. To ensure the visual-textual alignment, for text-rich images lacking text and language annotations, we subject them to a standardized data cleaning process, which includes multilingual text recognition and language classification. Afterwards, we organize all the text-rich images we have obtained into language-specific groups, preparing them for the subsequent stage of data annotation.
@@ 3.2 Human Expert Annotation
To obtain informative and accurate text-related QA pairs for images grouped by specific languages, a specialized group of annotators with expertise in the local regions of each language is recruited. These annotators are native speakers, guaranteeing a profound understanding and skillfulness in the linguistic subtleties and cultural contexts required for precise annotations. Given the subjective nature of understanding text within images, the annotation team is divided into two independent groups. One group is tasked with generating questions and providing answers based on the images, while the other group is responsible for evaluating and correcting the QA pairs. This division, known as the raise-then-correct paradigm, ensures a thorough and trustworthy evaluation of the text-rich image comprehension process. Moreover, the annotation results for each language underwent a 10% sampling inspection by a quality inspector to ensure adherence to standards. QA pairs that are not meeting the criteria are returned for re-annotation. Before beginning the formal annotation task, all annotators have received training and examples of annotations. The two-round annotation process is briefly illustrated in Fig. 4, further details of which are elaborated in the subsequent sections.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 5: Word clouds showcasing top answers in various languages, tokenized via NLTK with removing stop words, punctuation, and digits.
First Round Questioning and Answering. In the first round, we allocate three annotators per language to generate initial QA results. The annotators are tasked with thoroughly examining the textual and visual elements within a text-rich image in our collection. They are to examine the textual and visual elements within the image to formulate five meaningful and distinct questions and provide corresponding answers. All annotators should adhere to the following criteria: (1) the first three questions should satisfy that answering these questions requires direct reading of the textual information in the image, (2) the fourth and fifth questions require reasoning about the text in the image to answer, (3) the questions and answers must be reasonably correct and consistent with the content of the image, and (4) the answer should be as concise as possible and free of nonsense (e.g., when the question is “When is the volunteer recruitment period”, the answer should be “9:00-16:00” rather than “The volunteer recruitment period is 9:00-16:00”). This requirement for brevity aims to make the evaluation process user-friendly and reliable, avoiding the influence of extraneous content on the evaluation metrics.
Second Round Evaluation and Correction. To reduce the effect of human subjective cognitive bias on our MTVQA benchmark and get high-quality question-answer pairs, we assign two annotators for each language to carry out the annotation evaluation and correction process. These annotators follow a specific evaluation and correction protocol to ensure consistency and accuracy based on the first round results: (1) Assess the relevance of the question to the text in the image. Irrelevant QA pairs are discarded; (2) Verify the correctness of the answer and make necessary modifications; (3) Check for redundancy in the answer. If the answer repeats the question’s information, the repeated content is removed for conciseness. (4) Ethical Assessment. Review the image content or the question-answer pair for any unethical content, including but not limited to politics, personal privacy issues, etc. Such content is removed to uphold ethical standards in the dataset.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
3.3 Data Statistics
The MTVQA benchmark consists of 8,794 images and 28,607 question-answer pairs across the nine languages, divided into a training set with 6,678 images and 21,829 question-answer pairs and a test set with 2,116 images and 6,778 question-answer pairs. Detailed data distribution is illustrated in Fig. 2. Additionally, the benchmark showcases the vocabulary richness for each language through word clouds, as depicted in Fig. 5, and the lengths of questions and answers are statistically analyzed using the GPT-4o tokenizer, as shown in Fig. 6.
4 Experiments
4.1 Baseline Models
To comprehensively assess MLLMs’ multilingual perception and comprehension capabilities, We select state-of-the-art MLLMs from three categories: open-source general MLLMs, open-source text-centric MLLMs, and closed-source MLLMs. Each category contains the following models: (1) General MLLMs: InternVL-V1.5 [47], InternLM-Xcomposer2-4KHD [16], Mini-Gemini-HD-34B [17], Llava-Next-34B [18], DeepSeek-VL [8], YI-VL-34B [9]; (2) Text-centric MLLMs: TextSquare [14], TextMonkey [13], mPLUG-DocOwl 1.5 [12], MiniCPM-V 2.0 [48]; (3) Closed-source MLLMs: GPT-4o [1], GPT-4V [2], Gemini Ultra [4], QwenVL Max [7], QwenVL Plus [7], Claude3 Opus [5], Claude3 Sonnet [5], and GLM-4V [49]. For the closed-source MLLMs, we use the chat version through official APIs, while for the open-source MLLMs, we utilize the instruct versions available on the HuggingFace Model Hub. The open-source MLLMs’ model size varies from 7b to 34b.
4.2 Implementation Details
We conduct the evaluation experiments over the baseline MLLMs with their default settings, ignoring the effect of generation configuration on the results. To make the output of MLLMs more evaluation-friendly, we design the following prompt format to limit the output length: “Answer the question using a word or phrase in the language of the question. + <Question>”, where “<Question>” represents the actual question from the MTVQA test set. This approach aims to make the answers as concise as possible. Besides, InternLM-Xcomposer2-4KHD  [16] is chosen as the base model for an instruction tuning experiment on the MTVQA training set. The instruction tuning process adheres to the default training settings specified by the source, with “HD-16” and completes one epoch of training on 8 NVIDIA-A100 GPUs within 2 hours.
Table 1: Performance of the leading closed- and open-source MLLMs on the MTVQA benchmark. The best results of each language are bolded. The best results for open-source models are underlined. “Xcomposer-SFT” denotes instruction tuning to Xcomposer2-4KHD with MTVQA’s training set.
AR DE FR IT JA KO RU TH VI Avg.
Closed-Source MLLMs
GPT-4V [2] 11.5 31.5 40.4 32.3 11.5 16.7 10.3 15.0 28.9 22.0
GPT-4o [1] 20.2 34.2 41.2 32.7 20.0 33.9 11.5 22.5 34.2 27.8
Gemini Ultra [4] 14.7 32.3 40.0 31.8 12.3 17.2 11.8 20.3 28.6 23.2
QwenVL Max [7] 7.7 31.4 37.6 30.2 18.6 25.4 10.4 4.8 23.5 21.1
QwenVL Plus [7] 4.8 28.8 33.7 27.1 12.8 19.9 9.4 5.6 18.1 17.8
Claude3 Opus [5] 15.1 33.4 40.6 34.4 19.4 27.2 13.0 19.5 29.1 25.7
Claude3 Sonnet [5] 10.5 28.9 35.6 31.8 13.9 22.2 11.0 15.2 20.8 21.1
GLM-4V [49] 0.3 30.0 34.1 30.1 3.4 5.7 3.0 3.5 12.3 13.6
Open-Source MLLMs
InternVL-V1.5 [47] 3.4 27.1 31.4 27.1 9.9 9.0 4.9 8.7 12.4 14.9
Mini-Gemini-HD-34B [17] 2.2 25.0 29.2 25.5 6.1 8.6 4.1 4.3 11.8 13.0
Llava-Next-34B [18] 3.3 24.0 28.0 22.3 3.6 6.1 2.6 0.4 9.8 11.1
DeepSeek-VL [8] 0.6 14.2 15.3 15.2 2.9 3.8 1.6 0.9 5.2 6.6
YI-VL-34B [9] 1.7 13.5 15.7 12.1 4.8 5.2 0.8 3.5 4.1 6.8
MiniCPM-V 2.0 [48] 1.3 12.7 14.9 17.0 3.7 5.6 2.2 2.2 6.8 7.4
MiniCPM-V 2.5 [48] 6.1 29.6 35.7 26.0 12.1 13.1 5.3 12.6 15.3 17.3
TextSquare [14] 3.7 27.0 30.8 26.7 3.2 7.2 6.7 5.2 12.4 13.6
TextMonkey [13] 2.0 18.1 19.9 22.1 4.6 7.2 3.2 0.9 11.1 9.9
mPLUG-DocOwl 1.5 [12] 1.0 13.9 14.9 18.2 2.9 5.0 2.0 0.9 6.4 7.2
Xcomposer2-4KHD [16] 2.0 20.6 23.2 21.6 5.6 7.7 4.1 6.1 10.1 11.2
Xcomposer-SFT 11.8 31.7 37.4 29.3 14.5 12.9 5.8 13.9 20.2 19.7
4.3 Evaluation Results
Evaluation metric. To accurately assess whether the visual text that occurs in the answer is correct, we adopt Accuracy as the metric. The Accuracy metric measures the percentage of questions for which the predicted answer matches exactly with any of the target answers for the question. The accuracy metric awards a zero score even when the prediction differs slightly from the target answer.
[…]
Table 2: Few-shot performance of GPT-4V on the MTVQA benchmark. In-context examples are randomly selected from the training set of MTVQA in the respective languages. “n-shot” represents the number of the selected in-context examples.
AR DE FR IT JA KO RU TH VI Avg.
zero-shot 11.5 31.5 40.4 32.3 11.5 16.7 10.3 15.0 28.9 22.0
two-shot 11.6 35.3 42.2 33.1 13.2 17.0 11.4 19.0 31.0 23.7
five-shot 13.5 34.8 43.0 35.7 13.4 18.8 11.6 19.2 33.0 24.8
eight-shot 15.5 34.8 42.1 35.6 13.4 17.5 12.1 18.2 32.8 24.7
Few-shot evaluation of GPT-4V. Here, we compare the performance of GPT-4V on MTVQA under the few-shot settings. Specifically, we perform zero-shot, two-shot, five-shot, and eight-shot evaluations. We randomly select in-context examples from the train set in the respective languages and evaluate GPT-4V on the remaining instances. As shown in Tab. 2, compared to the zero-shot setting, GPT-4V’s performance has improved considerably under few-shot, highlighting its exceptional in-context learning ability in multilingual text comprehension contexts. Moreover, comparisons based on varying numbers of in-context samples reveal that augmenting the in-context samples can aid in further enhancement; however, after reaching a certain volume, the improvement becomes saturated.
5 Limitation
The current version of the MTVQA dataset, while dialectically diverse, exhibits limitations in its language coverage. Despite encompassing a range of languages, it falls short of inclusivity, omitting numerous lesser-spoken languages. This prompts our future continual endeavor to ensure comprehensive representation across the linguistic spectrum. Additionally, the dataset currently provides a single canonical response per question, which may not fully capture the range of answers corresponding to different expressions of the same underlying semantics. Recognizing the multifaceted nature of the inquiries, future versions will aim to include a spectrum of plausible answers to reflect the varied perspectives inherent to each question.
6 Conclusion
To attack the visual-textual misalignment issue in multilingual TEC-VQA, we introduce MTVQA, a new benchmark featuring high-quality human expert annotations in 9 diverse languages. We believe that MTVQA is the first benchmark to provide fully manual annotations tailored to text-centric scenarios. The results obtained from closed-source, general-purpose and text-centric MLLMs on our MTVQA dataset indicate that there is still room for improving their performance in multilingual text-centric scenarios. Although the MTVQA dataset has limitations, such as the underrepresentation of several lesser-spoken languages and single canonical answers, future updates will address these issues by expanding the multilingual scope and including a range of plausible answers. We are confident that this dataset can inspire researchers within the TEC-VQA community with new perspectives and ideas.
Appendix A Appendix
(a) German (DE) example
(b) Italian (IT) example
(c) French (FR) example
(d) Japanese (JA) example
(e) Korean (KO) example
Figure 7: Multilingual VQA examples selected from five languages. The corresponding translations in English are in brackets.